Flying against the wind, an airplane travels 2790 km in 3 hours. Flying with the wind, the same plane travels 10,160 km in 8 hours. What is the rate of the plane in still air and what is the rate of the wind?
Answer by richwmiller(17219)

3a=2790
a=2790/3=930
8w=10160
w=10160/8=1270
1270-930=340
340/2=170 speed of wind
930+170=1100 speed of plane
1270-170=1100 speed of plane